OpenStudy (anonymous):

one cylinder has a height of 16m the other cylinder has a height of 28m. If the smaller cylinder has a lateral area of approximately 302 square meters, what is the approximate lateral area of the larger cylinder? Answer 1,618.5 square meters 924.9 square meters 528.5 square meters 98.6 square meters

OpenStudy (mathteacher1729):

Are we assuming that the cylinders have the same radius?

OpenStudy (anonymous):

yes because it doesnt give the radius

OpenStudy (mathteacher1729):

Do you know the formula for lateral area of a cylinder?

OpenStudy (anonymous):

Lateral Area= (2pi)x(radius)x(height)

OpenStudy (mathteacher1729):

Yup. You can now solve for r by substituting the height and lateral area of the smaller cylinder of that formula. :)

OpenStudy (anonymous):

im still confused

OpenStudy (mathteacher1729):

Lateral Area= (2pi)x(radius)x(height) shorten this to LA = 2*pi*r*h You are given LA and h for the small cylinder. small cylinder LA = 302 m^2 small cylinder h = 16m Substitute those values into the formula and solve for r: LA = 2*pi*r*h 302 = 2*pi*r*16 302 = 32*pi*r (I multiplied 2 x 16 = 32) 302 / (32*pi) = r ( I divided both sides by 32*pi) 3.00404955 = r ( I entered 302 divided by (32 *pi) into a calculator, or even google!) So r = 3.00404955 Now we can solve for the LA of the large cylinder.

OpenStudy (anonymous):

so would it be 528.5

OpenStudy (mathteacher1729):

Yup! :D You got it.

OpenStudy (anonymous):

thank you :)

OpenStudy (mathteacher1729):

Hint: So you remember what you did later, it's best to actually write: LA = 2*pi*(3.00404955)*28 = 528 m^2

OpenStudy (anonymous):

your awesome

OpenStudy (mathteacher1729):

Aww thanks! :D